# A Simple Review of EEG Foundation Models: Datasets, Advancements and Future Perspectives

### v1

@@ Abstract
Electroencephalogram (EEG) signals play a crucial role in understanding brain activity and diagnosing neurological disorders. This review focuses on the recent development of EEG foundation models(EEG-FMs), which have shown great potential in processing and analyzing EEG data. We discuss various EEG-FMs, including their architectures, pre-training strategies, their pre-training and downstream datasets and other details. The review also highlights the challenges and future directions in this field, aiming to provide a comprehensive overview for researchers and practitioners interested in EEG analysis and related EEG-FMs.
Index Terms:
Electroencephalogram (EEG), EEG foundation model(EEG-FM).
@@ I Introduction
By addressing these essential aspects, this review paper will provide a comprehensive and in-depth analysis of the application of EEG-FMs for EEG tasks. The goal of this review is to provide a comprehensive overview of the current state-of-the-art EEG foundation models. The models we investigated include the following: Brant[19] [^1] Github: https://zju-brainnet.github.io/Brant.github.io/, BrainWave(Brant-2)[20] [^2] Github: https://github.com/yzz673/Brant-2, LaBraM[21] [^3] Github: https://github.com/935963004/LaBraM, NeuroGPT[22] [^4] Github: https://github.com/wenhui0206/NeuroGPT, BIOT[23] [^5] Github: https://github.com/ycq091044/BIOT, EEGPT1[24] [^6] Github: https://github.com/BINE022/EEGPT, BrainBERT[25] [^7] Github: https://github.com/czlwang/BrainBERT, FoME[26] [^8] Github: https://github.com/1061413241/FoME, EEGPT2[27], MBrain[28], NeuroLM[29], CBraMod[30] [^9] Github: https://github.com/wjq-learning/CBraMod and EEGFormer[31].
@@ II Overview of EEG Foundation Models
The key benefits of EEG-FMs over traditional deep learning approaches are their ability to handle large-scale, multichannel, and high-dimensional time-series data, as well as their capacity to generalize across different subjects and conditions. Like other large models, they are designed to scale with increasing data and computational resources, making them ideal for tasks requiring sophisticated understanding of brain activity. The challenge, however, lies in the computational demands of training such large models, which require significant GPU resources and sophisticated optimization techniques to ensure efficient learning. Additionally, there are ongoing efforts to improve model interpretability and avoid issues such as overfitting, where the model may fail to generalize to new or unseen EEG data. Recent advancements in techniques such as attention mechanisms and hybrid models that combine temporal and spatial processing are helping to address these challenges, further enhancing the potential of EEG-FMs.
@@ II-B1 Patching
Since neural recordings are electrical signals with high sampling rates, we aggregate timestamps into patches to (1) enhance the locality and extract semantic information; (2) reduce computation and memory usage; and (3) attend a longer temporal dependency[32]. Assume that the timestamp for each sample is $t$ and the stride is $s$. $X$ can be segmented into $\left\lfloor\frac{T-t}{s}+1\right\rfloor$, and each sample $\boldsymbol{x\in\mathbb{R}^{t\times C}}$, where $t$ is the number of timestamps and $C$ is the number of electrode channels, we divide $\boldsymbol{x}$ with length $M$, stride $S$ and overlap $O$, where $O=M-S$ is the number of overlapping timestamps between consecutive patches. This generate a set of patches $\boldsymbol{p\in\mathbb{R}^{N_{p}\times C\times M}}$, where $N_{p}=\left\lfloor\frac{t-M}{S}+1\right\rfloor$ is the number of patches in each channel. In fact, $\boldsymbol{p}$ can be represent as:
@@ II-B3 Frequency-Encoding
[…]
@@ II-C4 Codebook prediction
For Embedding prediction: We randomly generate a mask =$\mathcal{M}=\left\{{{m}_{i}\in{\left\{0,1\right\}}\mid i=1,\ldots,N_{p}}\right\}$. After that, we replace the masked patches of $p$ with the learnable mask token $e_{\mathcal{M}}\in{\mathbb{R}^{D}}$. The corrupted EEG patches can be denoted as $e^{\mathcal{M}}=\left\{e_{i}:m_{i}=0|i=1,2,\dots,N_{p}\right\}\cup\left\{e_{\mathcal{M}}:m_{i}=1|i=1,2,\dots,N_{p}\right\}$. , which will be added by subsequent modules. $\boldsymbol{\bar{p}}\in\mathbb{R}^{N_{p}\times C\times D}$ are used to predict the corresponding neural tokens through a linear classifier:
@@ III-A Pretraining and Downstream Datasets
In addition, to evaluate the performance of EEG foundation models in downstream tasks, it is necessary not only to use standard performance indicators, but also to pay attention to the model’s generalization ability, interpretability, robustness, computational efficiency and other dimensions. Therefore, the selection of downstream task data also needs to take into account the relevance and complexity of the task. We have collected all relevant datasets for pre-training and downstream tasks for all the models, see the Tab.LABEL:tab:Datasets_for_pretraining_and_downstream_tasks and Tab.LABEL:tab:Datasets_for_pretraining_and_downstream_tasks(continued) .
@@
  • 1
  • 2
    * Indicates the corresponding dataset is private .
  • 3
TABLE III: Datasets for pretraining and downstream tasks(continued)
Pretraining Datasets Downstream Datasets
BrainBERT[25] - -
@@ III-B Datasets preprocessing
TABLE IV: The preprocessing conditions of EEG datasets
@@ III-C Parameter Size
TABLE V: The parameter size of EEG-FMs.
@@ III-D Downstream Task Types
TABLE VI: The types of downstream tasks
@@ III-E Self-Supervise Learning
TABLE VII: The reconstruction objectives of EEG-FM
@@ III-F Hardware for pretraining
TABLE VIII: The hardware conditions for the pre-training of EEG-FM
@@ IV Discussion
IV-A Model Architectures
Approaches such as transformer-based frameworks (e.g., EEGPT and BrainBERT) have demonstrated their efficacy in handling the temporal and spatial complexities of EEG data. For instance, EEGPT’s electrode-wise modeling and hierarchical design allow for independent processing of spatial and temporal features, which is crucial given the high variability and noise inherent in EEG signals. Meanwhile, FoME’s adaptive temporal-lateral attention scaling enhances its ability to model long-range dependencies in EEG recordings. Models like Brant-2 and LaBraM have integrated multimodal capabilities by leveraging shared representations across invasive and non-invasive signals, further advancing the adaptability and robustness of EEG models.
IV-B Pre-training Strategies and Data Utilization
Pre-training strategies have been pivotal in overcoming the scarcity of labeled EEG data. Self-supervised learning, including masked autoencoding and autoregressive modeling, has emerged as a cornerstone for many EEG foundation models. For example, EEGPT and Brant-2 utilize masked pretraining to learn robust and generalizable representations, enabling effective fine-tuning across downstream tasks. The incorporation of massive datasets, such as those used in LaBraM and Brant, has also contributed significantly to the generalizability of these models. However, these approaches often require meticulous preprocessing and alignment of diverse datasets, which remains a logistical and computational challenge.
IV-C Performance Evaluation
EEG foundation models have been extensively evaluated across a wide range of tasks, including emotion recognition, cognitive state classification, seizure detection, and motor imagery. However, the lack of standardized benchmarks and datasets hinders direct comparisons and may obscure the full capabilities of these models.
IV-D Future Directions
@@
Secondly, GNN can be considered to capture the spatial pattern of EEG signals. We found that EEGPT2 [27] and [103] has tried to integrate GNN into EEG-FM to better capture the potential spatial features of EEG signals. Thirdly, the number of parameters of EEG-FM is still small compared to the current mainstream large language models(e.g. ChatGPT-4o). How to design models that fit specific tasks well but have small parameters, and how to design models with large parameters and strong generalization will be the focus of future work. Fourthly, designing a model that generates EEG signals from language will be a competitive direction, which may solve the problem of lack of model pre-training EEG signals. Finally, combining EEG models with other modes such as language and imaging to create multimodal vertical fields EEG-FM is expected to create an overall framework connecting neuroscience and artificial intelligence, opening up new avenues for research and application.

### v2

@@ Abstract
Electroencephalogram (EEG) signals play a crucial role in understanding brain activity and diagnosing neurological diseases. Because supervised EEG encoders are unable to learn robust EEG patterns and rely too heavily on expensive signal annotation, research has turned to general-purpose self-supervised EEG encoders, known as EEG-based models (EEG-FMs), to achieve robust and scalable EEG feature extraction. However, the readiness of early EEG-FMs for practical applications and the standards for long-term research progress remain unclear. Therefore, a systematic and comprehensive review of first-generation EEG-FMs is necessary to understand their current state-of-the-art and identify key directions for future EEG-FMs. To this end, this study reviews 14 early EEG-FMs and provides a critical comprehensive analysis of their methodologies, empirical findings, and unaddressed research gaps. This review focuses on the latest developments in EEG-based models (EEG-FMs), which have shown great potential for processing and analyzing EEG data. We discuss various EEG-FMs, including their architectures, pretraining strategies, pretraining and downstream datasets, and other details. This review also highlights challenges and future directions in the field, aiming to provide a comprehensive overview for researchers and practitioners interested in EEG analysis and related EEG-FM.
@@ I Introduction
By addressing these essential aspects, this review paper will provide a comprehensive and in-depth analysis of the application of EEG-FMs for EEG tasks. The goal of this review is to provide a comprehensive overview of the current state-of-the-art EEG foundation models. The models we investigated include the following: Brant[19] [^1] https://zju-brainnet.github.io/Brant.github.io/, BrainWave(Brant-2)[20] [^2] https://github.com/yzz673/Brant-2, LaBraM[21] [^3] https://github.com/935963004/LaBraM, NeuroGPT[22] [^4] https://github.com/wenhui0206/NeuroGPT, BIOT[23] [^5] https://github.com/ycq091044/BIOT, EEGPT1[24] [^6] https://github.com/BINE022/EEGPT, BrainBERT[25] [^7] https://github.com/czlwang/BrainBERT, FoME[26] [^8] https://github.com/1061413241/FoME, EEGPT2[27], MBrain[28], NeuroLM[29], CBraMod[30] [^9] https://github.com/wjq-learning/CBraMod, EEGFormer[31] and ALFEE[32] [^10] https://github.com/xw1216/ALFEE.
@@ II-B1 Patching
Since neural recordings are electrical signals with high sampling rates, we aggregate timestamps into patches to (1) enhance the locality and extract semantic information; (2) reduce computation and memory usage; and (3) attend a longer temporal dependency[33]. Assume that the timestamp for each sample is $t$ and the stride is $s$. $X$ can be segmented into $\left\lfloor\frac{T-t}{s}+1\right\rfloor$, and each sample $\boldsymbol{x\in\mathbb{R}^{t\times C}}$, where $t$ is the number of timestamps and $C$ is the number of electrode channels, we divide $\boldsymbol{x}$ with length $M$, stride $S$ and overlap $O$, where O=MSO=M-S is the number of overlapping timestamps between consecutive patches. This generate a set of patches $\boldsymbol{p\in\mathbb{R}^{N_{p}\times C\times M}}$, where $N_{p}=\left\lfloor\frac{t-M}{S}+1\right\rfloor$ is the number of patches in each channel. In fact, $\boldsymbol{p}$ can be represent as:
@@ II-B3 Frequency-Encoding
$P_{k,c}(f)=\frac{1}{T}\left|\mathcal{F}\{P_{k,c}(t)\}\right|^{2}=\frac{1}{T}\left|\int_{-\infty}^{\infty}P_{k,c}(t)e^{-i2\pi ft}dt\right|^{2}$ (2)
@@ II-C4 Codebook prediction
For Embedding prediction: We randomly generate a mask $\mathcal{M}=\left\{{{m}_{i}\in{\left\{0,1\right\}}\mid i=1,\ldots,N_{p}}\right\}$. After that, we replace the masked patches of $p$ with the learnable mask token $e_{\mathcal{M}}\in{\mathbb{R}^{D}}$. The corrupted EEG patches can be denoted as $e^{\mathcal{M}}=\left\{e_{i}:m_{i}=0|i=1,2,\dots,N_{p}\right\}\cup\left\{e_{\mathcal{M}}:m_{i}=1|i=1,2,\dots,N_{p}\right\}$. , which will be added by subsequent modules. $\boldsymbol{\bar{p}}\in\mathbb{R}^{N_{p}\times C\times D}$ are used to predict the corresponding neural tokens through a linear classifier:
@@ III-A Pretraining and Downstream Datasets
In addition, to evaluate the performance of EEG foundation models in downstream tasks, it is necessary not only to use standard performance indicators, but also to pay attention to the model’s generalization ability, interpretability, robustness, computational efficiency and other dimensions. Therefore, the selection of downstream task data also needs to take into account the relevance and complexity of the task. We have collected all relevant datasets for pre-training and downstream tasks for all the models, see the Tab.LABEL:tab:Datasets_for_pretraining_and_downstream_tasks Tab.LABEL:tab:Datasets_for_pretraining_and_downstream_tasks(continued1) and Tab.LABEL:tab:Datasets_for_pretraining_and_downstream_tasks(continued2). As can be seen from the table, different EEG-FMs utilize varying pre-training datasets and downstream task datasets, which can pose challenges. The size and quality of pre-training data can impact EEG-FM performance, making it difficult to objectively compare the best performance of different EEG-FMs on downstream tasks.
@@
  • 1
  • 2
    * Indicates the corresponding dataset is private .
  • 3
TABLE III: Datasets for pretraining and downstream tasks(continued1)
Pretraining Datasets Downstream Datasets
@@
  • 1
  • 2
    * Indicates the corresponding dataset is private .
  • 3
TABLE IV: Datasets for pretraining and downstream tasks(continued2)
Pretraining Datasets Downstream Datasets
BrainBERT[25] - -
@@
ALFEE[32] TUEG[46] TUAB[57]
SEED-IV[71] TUAR[67]
SEED-V[60] TUSL[69]
SEED-GER[71] TUSZ[57]
SEED-FRA[91] Neonate[103]
BCIC4-1[56] TUAB[57]
Emobrain[58] TUEV[57]
Grasp and Lift[59] SEED[70]
Inria BCI[61] HMC[96]
MMI[63] TUSL[69]
RSEEG[65] Workload[94]
RawEEG[104] -
Siena[43] -
SPIS[99] -
TVNT[66] -
@@ III-B Datasets preprocessing
Most EEG-FMs perform minimal and simple data preprocessing steps, namely filtering and resampling, to standardize EEGs from various sources. Notably, the removal of noise-related outliers, suppression of EEG artifacts, artifactual EEG components, and site-related harmonization were not explicitly pursued. Moreover, data normalization strategies that produce training-ready samples were not sufficiently described in most studies. It remains unclear whether or how various offline data handling strategies impact EEG-FM pretraining and downstream task performance.
TABLE V: The preprocessing conditions of EEG datasets
@@ III-C Parameter Size
TABLE VI: The parameter size of EEG-FMs.
@@
ALFEE[32] ALFEE-S 16.3M
ALFEE-M 44.3M
ALFEE-B 120M
ALFEE-L 300M
ALFEE-XL 540M
@@
Aside from ALFEE[32], few EEG-FM related papers have examined and analyzed the impact of model and data scaling on task performance. Regarding model scaling, most papers lack comprehensive experimental evidence, making it unclear whether a clear trend in model scaling exists. Regarding data scaling, some papers conclude that increasing preprocessed data may lead to poorer performance on downstream tasks or the evidence for data scaling is weak. ALFEE[32] demonstrates that both model and data scaling apply equally to EEG-FM, with larger models and more preprocessed data generally resulting in better performance.
@@ III-D Downstream Task Types
TABLE VII: The types of downstream tasks
@@
ALFEE[32] Abnormal detection $\times$ $\times$ $\times$ -
Event Type Classification
Emotion Recognition
Sleep Stage Classification
Workload Detection
Slowing Event Detection
@@ III-E Self-Supervise Learning
TABLE VIII: The reconstruction objectives of EEG-FM
@@
ALFEE[32] EEG signal forecasting see [32]
Original patches
Task classification
@@ III-F Hardware for pretraining
TABLE IX: The hardware conditions for the pre-training of EEG-FM
@@ IV Discussion
IV-A Current Research Challenges
Effects of Preprocessing and Normalization: EEG datasets may require extensive offline preprocessing to manage data quality and suppress artifacts. Because the current study performed minimal processing, it is unclear whether data outliers and noise impact EEG-FM preprocessing and task performance. Even clean EEG datasets require careful consideration of normalization strategies, as EEG can vary across channels, subjects, and acquisition sites[105]. Therefore, there is a need to better understand the impact of these choices on downstream EEG-FM modeling.
Standardized Benchmark Tasks: The lack of common tasks in EEG-FM evaluation makes understanding the state of the art challenging and highlights the need to identify a common core evaluation set for future EEG-FM evaluation. This set must cover a variety of task types, including classification and regression, and have both dense (one label for the entire EEG recording) and sparse (one label for each EEG segment) labels. Furthermore, some existing datasets may have saturated the performance of traditional methods[106], preventing them from fully exploring the potential of EEG-FM. Therefore, collecting and open-sourcing challenging new datasets is crucial for the future development of EEG-FM.
Long-Term Context Modeling: The required EEG recording duration depends on the signal-to-noise ratio of the target neural signal, the statistical requirements for improving the signal-to-noise ratio through repeated sampling, and the temporal characteristics of the brain activity being studied.These include:
  • Transient events (milliseconds to seconds): ERPs[107, 108, 109, 110], such as the P300 (peak approximately 300 ms after stimulus), the N170 (peak approximately 170 ms), and the error-related negativity (ERN, peak approximately 100 ms after the response), are brief, rapid responses to discrete stimuli or events. The ”unit of analysis” here is a short time window (epoch) surrounding the event, typically 1-2 seconds.
  • Rhythmic states (seconds to minutes): Resting-state or task-state[111] EEG analysis focuses on persistent EEG oscillations that remain relatively stable over several minutes, such as alpha, beta and specific task-related rhythms. In these cases, the ”unit of analysis” is a continuous recording lasting several minutes.
  • Evolutionary states (tens of minutes to hours): Cognitive states[112, 113, 114, 115], such as driving fatigue and mental workload, evolve gradually during task performance. The ”unit of analysis” here is the duration of the entire task, as the study is precisely about changes in states over time.
  • Macrostructural Patterns (Hours): Sleep staging[116, 117] aims to observe the complete structure of sleep, including multiple approximately 90-minute cycles of non-rapid eye movement (NREM) and rapid eye movement (REM) sleep. This requires a full night of recording, typically 6-8 hours. The ”unit of analysis” is the entire sleep period.
Trustworthy Modeling: Currently, no EEG-FMs focus on model interpretability or explanation, which remains a key requirement for data-driven modeling in high-stakes and expert-centric medical fields. Research is needed to uncover the black box of EEG-FMs and gain insights into the actual robustness of their decision-making processes, both in pre-training and in downstream applications. Linking to known brain physiological or pathological patterns may be necessary for expert users to better assist in EEG-based diagnosis of disorders (such as ASD, ADHD, and MDD).
IV-B Performance Evaluation
One of the most significant obstacles in current EEG-FM research is the lack of standardized evaluation protocols and benchmarks[118]. Different studies evaluate their models on a wide range of downstream tasks and datasets, making direct and fair comparisons nearly impossible. For example, one model may perform well on a seizure detection task, while another performs poorly on an emotion recognition task. Even for the same task (e.g., motor imagery), the datasets, preprocessing pipelines, and data partitioning strategies (e.g., within-subject vs. between-subject cross-validation) used can vary, making performance results incomparable. This heterogeneity hinders a clear understanding of true progress in the field. To address this issue, the research community urgently needs to develop and adopt standardized benchmark suites, such as emerging initiatives like AdaBrain-Bench[119] [^11] https://github.com/Jamine-W/AdaBrain-Bench, which aims to provide a framework covering a variety of BCI tasks, multi-dimensional evaluation metrics, and standardized adaptation procedures. Such benchmarks will provide a level playing field for evaluating generalization across subjects, tasks, and datasets.
Many models are evaluated under flawed settings. A common problem is ”in-sample” evaluation, where the downstream task data used for fine-tuning and testing is itself part of the pre-training dataset. This approach fails to truly test the model’s generalization ability to new, unseen data. A more rigorous and realistic evaluation metric is ”out-of-distribution” (OOD) evaluation, where model performance is tested on datasets completely unseen during model pre-training. This simulates real-world scenarios when deploying the model to a new hospital or using it with new hardware. Furthermore, the evaluation should clearly distinguish between cross-subject generalization (i.e., performance on new subjects in the same dataset) and the more challenging cross-dataset generalization (i.e., performance on datasets from different studies with different protocols and hardware). Only through rigorous OOD and cross-dataset testing can we truly trust that the model has learned universal neurophysiological representations, rather than biases or artifacts from a specific dataset.
In addition, methods for evaluating EEG-FM capabilities also need to diversify. Currently, the dominant evaluation paradigm is to fully fine-tune the entire pre-trained model on a downstream task. While this demonstrates the model’s best potential performance, it does not fully reveal the intrinsic quality of the representations learned during pre-training[120].
  • Linear probing: In this strategy, the pre-trained encoder weights are frozen, and only a simple linear classifier is trained on top of them. The performance of linear probing directly reflects the quality and separability of the learned representations. However, many studies have reported that linear probing performs significantly worse than full fine-tuning, raising questions about the intrinsic quality of representations learned by current self-supervised learning strategies.
  • Few-Shot and Zero-Shot Learning: This is a critical test of the capabilities of a base model. Evaluating a model’s ability to handle new tasks with only minimal (few-shot) or even no (zero-shot) labeled training examples is central to measuring its true utility as a general-purpose feature extractor.
Finally, evaluation metrics themselves need to go beyond traditional classification accuracy to better reflect the model’s actual clinical or application value. For example, in epilepsy detection tasks, false positives per hour is a more clinically relevant metric than the F1 score because it directly relates to alarm fatigue and clinical workflow[121]. Similarly, when assessing continuous states such as stress or fatigue, models must not only be accurate but also capable of early prediction[122], which requires specialized evaluation metrics. Therefore, future evaluation frameworks should integrate more application-specific and meaningful metrics to comprehensively assess a model’s real-world utility.
IV-C Future Directions
@@
Secondly, GNN can be considered to capture the spatial pattern of EEG signals. We found that EEGPT2 [27] and [123] has tried to integrate GNN into EEG-FM to better capture the potential spatial features of EEG signals. Thirdly, the number of parameters of EEG-FM is still small compared to the current mainstream large language models(e.g. ChatGPT-4o). How to design models that fit specific tasks well but have small parameters, and how to design models with large parameters and strong generalization will be the focus of future work. Fourthly, designing a model that generates EEG signals from language will be a competitive direction, which may solve the problem of lack of model pre-training EEG signals. Finally, current work primarily focuses on robust EEG representation, further research is needed to extend the model to multimodal scenarios such as Video-EEG, Image-EEG, and Text-EEG. Therefore, combining EEG models with other modes such as language and imaging to create multimodal vertical fields EEG-FM is expected to create an overall framework connecting neuroscience and artificial intelligence, opening up new avenues for research and application.